# OpenGraph: Towards Open Graph Foundation Models

### v2

@@ 3.1.2. Topology-aware Projection with Arbitary Graphs
To handle the significant variability in the dimensions of adjacent matrix $\tilde{\textbf{A}}$ across different graph datasets, using a fixed-input-dimension random neural network is not feasible. To address this challenge, we propose a topology-aware projection approach using a function $\phi:\mathbb{R}^{|\mathcal{V}|}\rightarrow\mathbb{R}^{d}$. By eliminating one of the varying dimensions, we aim to reduce information loss during dimension reduction. To achieve this, we employ a large hidden dimensionality $d$. Prior research (shen2021powerful, ; zheng2022instant, ) has demonstrated that even random projections can yield satisfactory performance when utilizing a large dimensionality. Hence, to better preserve graph information and minimize randomness in our graph tokenizer, we utilize the fast singular value decomposition (SVD) for the projection function $\phi$. Empirical analysis reveals that two iterations of fast SVD effectively preserve topology information while incurring negligible computational overhead. Formally, our graph tokenizer calculates the resulting token sequence based on the following operations:
@@
(4) $\textbf{e}_{v}=\phi(\tilde{\textbf{A}}_{v,:})=\tilde{\textbf{A}}_{v,:}\text{LN}((\textbf{U}\sqrt{\Lambda}\leavevmode\nobreak\ \|\leavevmode\nobreak\ \textbf{V}\sqrt{\Lambda})),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \textbf{U},\Lambda,\textbf{V}=\text{SVD}(\bar{\textbf{A}})$
@@
It is worth emphasizing that our graph tokenizer is designed to handle arbitrary variable-sized graphs that undergo significant shifts in terms of node token sets and relational semantics.
3.2. Scalable Graph Transformer
@@
After acquiring the universal topology-aware tokens for all nodes, regardless of any variations in characteristics across diverse graphs, through our unified graph tokenizer, the subsequent task is to empower our graph model to grasp the complex node-wise dependencies within the global context. Drawing inspiration from the success of transformer architectures in modeling complex relationships between instances, our OpenGraph utilizes a graph transformer as the backbone to incorporate global dependencies among all nodes.
To ensure scalability and effectively handle large-scale graph datasets, we introduce the following techniques, to enhance the efficiency of our graph transformer backbone in our OpenGraph.
3.2.1. Token Sequence Sampling
@@
To optimize the efficiency of our model, given the large embedding size, our graph transformer is trained using sampled token sequences from the current training batch. Specifically, the training batch $\{(v_{c_{b}},v_{p_{b}},v_{n_{b}})|b=1,\cdots,B\}$ consists of $B$ triplets of centric nodes $v_{c_{b}}$, positive nodes $v_{p_{b}}$, and negative nodes $v_{n_{b}}$, which serve as input for our graph transformer.
(5) $(\textbf{e}_{c_{1}},\textbf{e}_{c_{2}},\cdots\textbf{e}_{c_{B}})\leavevmode\nobreak\ \|\leavevmode\nobreak\ (\textbf{e}_{p_{1}},\textbf{e}_{p_{2}},\cdots,\textbf{e}_{p_{B}})\leavevmode\nobreak\ \|\leavevmode\nobreak\ (\textbf{e}_{n_{1}},\textbf{e}_{n_{2}},\cdots,\textbf{e}_{n_{B}})$
The initial token embeddings ${\textbf{e}_{v}}$ effectively capture the local structural information for each node $v\in\mathcal{V}$. By using the sampled token sequence, we can preserve the graph context with a particular emphasis on the training batch. This approach significantly reduces the sequence length from $|\mathcal{V}|$ to $3\times B$, enabling efficient training for large-scale graphs. Additionally, since the initial token embeddings $\textbf{e}_{v}$ inherently encode the topological closeness among nodes, our graph transformer does not rely on temporal embeddings typically used in transformers for temporal or textual sequences.
3.2.2. Efficient Self-attention with Anchor Sampling
Though sequence sampling reduces the length of the node sequence, the pairwise relation learning in self-attention still exhibits a complexity of $\mathcal{O}(B^{2}\times d)$, which can limit the batch size $B$ required for computational efficiency. To address this challenge, we introduce an additional step of sampling a set of anchor nodes ${v_{a_{s}}}$, where $s=1,\cdots,S$, during the self-attention calculation. Here, $S$ is chosen to be $d/H<B$ to ensure that the self-attention module incurs similar memory costs as other fully-connected components. Here, $H$ represents the number of attention heads. More specifically, the self-attention process for each head can be summarized as follows:
$\textbf{e}_{t}^{(3)}$ $=\sum_{v_{a}}\alpha_{t,a}\textbf{W}^{(v)}\textbf{e}_{a}^{(2)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \textbf{e}_{a}^{(2)}=\sum_{v_{t}}\alpha_{a,t}\textbf{W}^{(v)}\textbf{e}_{t}^{(1)}$
(6) $\alpha_{t,a}$ $=\text{softmax}\Big{(}(\textbf{W}^{(q)}\textbf{e}_{t})^{\top}\cdot(\textbf{W}^{(k)}\textbf{e}_{a})\leavevmode\nobreak\ /\leavevmode\nobreak\ \sqrt{d/H}\Big{)}$
Our efficient self-attention involves embeddings $\textbf{e}_{*}^{(1)},\textbf{e}_{*}^{(2)},\textbf{e}_{*}^{(3)}$ for anchor nodes $v_{a}$ and vanilla nodes $v_{t}$. After each attention calculation, the results from multiple heads are concatenated, passed through a learnable linear layer, and connected with a residual connection. The parameters $\textbf{W}^{(q)},\textbf{W}^{(k)},\textbf{W}^{(v)}$ are the parameters of the attention layer. To reduce computational complexity, we employ a two-stage self-attention process. It transforms the $3\times B$-length sequence to a shorter $S$-length sequence and then reverses the process. This decomposition reduces the complexity from $\mathcal{O}(B^{2}\times d)$ to $\mathcal{O}(B\times S)$, ensuring scalability for large-scale models.
After the self-attention module, each layer of our scalable graph transformer includes a two-layer fully-connected block with residual connections, accompanied by two layer normalization modules. To ensure numerical stability, per-layer scaling is applied. For a more comprehensive understanding of the specific configurations of our scalable graph transformer, please refer to Appendix A.
3.3. Knowledge Distillation from LLM
@@
Obtaining diverse graph datasets for different domains can be challenging due to factors like privacy issues that restrict access to essential data (zheleva2007preserving, ). Inspired by the remarkable knowledge and understanding demonstrated by large language models (LLMs), we leverage their power to enhance the generation of diverse graph-structured data. To improve the efficacy of our LLM-augmented graph data for pre-training our model, we have developed an augmentation mechanism. This mechanism enables the LLM-augmented graph data to closely approximate real-world graph characteristics, enhancing the relevance and usefulness of the augmented data.
3.3.1. LLM-based Node Generation
@@
When generating graphs, our initial step is to create a node set tailored to the specific application scenario. Each node is characterized by a text-based profile that assists in generating subsequent edges. However, this task can be particularly challenging when dealing with real-world scenarios due to the substantial scale of the node set. For instance, in e-commerce platforms, the graph data may consist of billions of products. As a result, efficiently enabling the LLM to generate a large number of nodes becomes a significant challenge.
@@
To address the above challenge, we employ a strategy that involves iteratively dividing general nodes into sub-categories with finer semantic granularity. For example, when generating product nodes, we begin by prompting the LLM with a query like "List all sub-categories of products on an e-commerce platform like Amazon." The LLM responds with a list of sub-categories such as "clothing," "home & kitchen," "electronics," and so on. We then continue this iterative division process by asking the LLM to further refine each sub-category. This process is repeated until we obtain nodes that closely resemble real-world instances, such as a product with labels like "clothing," "women’s clothing," "sweaters," "hooded sweaters," "white hooded sweaters." Appendix A.3.1 provides more details about our prompt template and specific generation examples.
@@
3.3.2. LLM-based Edge Generation by Gibbs Sampling
@@
To generate edges, we use the Gibbs sampling algorithm (gelfand2000gibbs, ) with the generated node set $\mathcal{V}$. The algorithm starts with a random sample, which varies depending on the type of entity-entity relation data. For instance, in a paper-wise citation network, the sample is a node pair $(v_{s_{0}},v_{t_{0}})$. In a person-entity relation scenario like an author-paper or user-item recommendation network, the initial sample is a binary vector $\textbf{a}^{0}\in\{0,1\}^{|\mathcal{V}|}$. Each element $a_{i}=1$ indicates interaction between the sampled person and the $i$-th node $v_{i}$, while $a_{i}=0$ indicates no interaction. In the case of person-entity relations, the Gibbs algorithm for edge sampling is described in Appendix A.3.3. The key is estimating the probability $p(\mathbf{a}^{t}\oplus v_{t^{\prime}}|\mathbf{a}^{t})$, with $\oplus$ representing setting the $t^{\prime}$-th dimension of $\textbf{a}^{t}$ to 1.
@@
Node-wise Connection Probability Estimation. To accurately estimate the probability $p(\mathbf{a}^{t}\oplus v_{t^{\prime}}|\mathbf{a}^{t})$ of connecting two nodes in our generated graph, we leverage the reasoning capabilities of the LLM. However, directly prompting the LLM for predictions on each edge can be computationally expensive, with the number of required prompts scaling proportionally to the number of edges ($\mathcal{O}(|\mathcal{V}|\times|\mathcal{V}|)$). To ensure efficient probability estimation, we adopt an alternative approach. We prompt the LLM to generate hidden representations $\mathbf{h}_{i}$ for each node $v_{i}$. Then, we calculate the relationship between each edge using dot-product similarity as:
(7) $p(\textbf{a}^{t}\oplus v_{t^{\prime}}|\textbf{a}^{t})=\sum_{v_{i}}a^{t}_{i}(\textbf{h}_{i}/\|\textbf{a}^{t}\|_{0})^{\top}\cdot\textbf{h}_{t^{\prime}}$
@@
By utilizing the text embeddings $\textbf{h}_{i}$ and $\textbf{h}_{t^{\prime}}$ provided by the LLM, which are in the same representation space, we can effectively capture the semantic relations between the respective nodes.
@@
Dynamic Probability Normalization. To ensure that the calculated probability scores fall within the range of $[0,1]$, our generation algorithm incorporates a dynamic probability normalization approach, which enhances the algorithm’s realism by mimicking real-world distributions. It maintains a record of the most recent $T^{\prime}$ estimation values, denoted as $\mathcal{P}=\{p(\textbf{a}^{t}\oplus v_{t^{\prime}}|\textbf{a}^{t})\leavevmode\nobreak\ |\leavevmode\nobreak\ t=-1,-2,\cdots,-T^{\prime}\}$. By calculating the mean ($\mu$) and standard deviation ($\sigma$) of these values, we obtain an understanding of their distribution. To adjust new estimations, we use $\mu+2\sigma$ as the upper bound and $\mu-2\sigma$ as the lower bound. The resulting adjusted probability score, denoted as $\bar{p}$, is obtained by $\bar{p}=(p-\mu)/(4\sigma)$. Through this normalization process, we ensure that the probability scores are constrained within a range that closely approximates $[0,1]$.
@@
Node Locality Incorporation. Our LLM-based edge generation method effectively determines potential connections based on the semantic similarity between nodes. However, it tends to create excessive connections among all semantically-related nodes, overlooking the important concept of locality observed in real-world graphs. In reality, nodes are more likely to be connected to a subset of relevant nodes, as they have limited interactions with only a portion of all other nodes. To address this, we introduce a way to incorporate the notion of locality in our edge generation process. Each node $v_{i}$ is randomly assigned a locality index $n_{i}$. To account for the difference in locality between two nodes, we use the absolute difference $|n_{i}-n_{j}|$ in an exponential decay function applied to the normalized probability. This results in an adjusted probability $\hat{p}$ calculated as $\hat{p}=\bar{p}\cdot\alpha^{|n_{i}-n_{j}|}$, where $10<\alpha<$ 1.
3.3.3. Graph Topological Pattern Injection
@@
4. Evaluation
Our experiments aim to address the following research questions:
  • RQ1: How does OpenGraph compare to existing graph models in zero-shot link prediction and node classification tasks?
  • RQ2: How does different training datasets (real data and generated data) affect the model performance of OpenGraph?
  • RQ3: What is the influence of different types of initial graph projection methods on our OpenGraph framework?
  • RQ4: How does the sampling strategies in our scalable graph transformer impact the model efficiency and performance?
  • RQ5: How does model scale impact our OpenGraph?
  • RQ6: How does our OpenGraph perform relative to baseline methods in standard full-shot training and few-shot training?
4.1. Experimental Settings
4.1.1. Datasets
We evaluate our OpenGraph on two graph learning tasks: link prediction and node classification, using a total of 8 real-world datasets. Appendix A.2.1 provides detailed descriptions.
4.1.2. Evaluation Protocol
Following previous works (he2020lightgcn, ; kipf2016semi, ), we adopt the original train-test data split for the experimental datasets. We pre-train our OpenGraph on generated datasets and conducts zero-shot prediction for the evaluation datasets made of real graph data. As most baselines struggle with cross-dataset transferring, we evaluate them in two few-shot training settings. Please refer to Appendix A.2.2 for more details about our cross-dataset zero-shot setting, few-shot settings, and evaluation metrics.
4.1.3. Implementation Details
We provide detailed information about the implementation of OpenGraph and the baseline methods, as well as the graph generation process, in Appendix A.2.3.
Table 1. Performance comparison between our OpenGraph (zero-shot) and baseline methods (one-shot, five-shot) on the link prediction task (measured by Recall@N for $N=20,40$) and node classification task (measured by Accuracy and Macro F1 Score).
@@ 4.1.3. Implementation Details
Metric R@20 R@40 R@20 R@40 R@20 R@40 R@20 R@40 R@20 R@40 Acc MacF1 Acc MacF1 Acc MacF1
MF 1-shot 0.0087 0.0161 0.0261 0.0349 0.0331 0.0604 0.1396 0.1956 0.0034 0.0043 0.1710 0.1563 0.1740 0.1727 0.3470 0.3346
5-shot 0.0536 0.0884 0.0412 0.0609 0.0987 0.1584 0.2060 0.2989 0.0196 0.0284 0.1500 0.1422 0.1520 0.1484 0.3540 0.3435
MLP 1-shot 0.0195 0.0336 0.0112 0.0185 0.0548 0.1019 0.1492 0.2048 0.0017 0.0028 0.2300 0.1100 0.2590 0.1993 0.4430 0.3114
5-shot 0.0621 0.1038 0.0115 0.0185 0.0851 0.1470 0.2362 0.2563 0.0092 0.0152 0.3930 0.3367 0.3690 0.3032 0.5240 0.4767
GCN 1-shot 0.0279 0.0459 0.0206 0.0321 0.0432 0.0849 0.1760 0.2086 0.0096 0.0160 0.3180 0.1643 0.3200 0.2096 0.4270 0.3296
5-shot 0.0705 0.1312 0.0366 0.0513 0.1054 0.1656 0.2127 0.2324 0.0251 0.0408 0.5470 0.5008 0.4910 0.4190 0.509 0.4455
GAT 1-shot 0.0580 0.1061 0.0258 0.0372 0.0245 0.0520 0.1615 0.2476 0.0047 0.0079 0.2420 0.1687 0.2810 0.2025 0.4720 0.3657
5-shot 0.0711 0.1309 0.0340 0.0505 0.1506 0.2267 0.2002 0.2883 0.0228 0.0392 0.585 0.5438 0.4940 0.4441 0.5780 0.5582
GIN 1-shot 0.0530 0.1004 0.0163 0.0247 0.0466 0.0884 0.1541 0.2388 0.0069 0.0114 0.3190 0.1753 0.2820 0.1705 0.4410 0.3064
5-shot 0.0735 0.1441 0.0311 0.0458 0.1458 0.2344 0.1926 0.2829 0.0252 0.0418 0.5400 0.4941 0.521 0.4696 0.5070 0.4547
DGI 1-shot 0.0315 0.0617 0.0255 0.0385 0.0486 0.0863 0.1868 0.2716 0.0081 0.0142 0.3150 0.1782 0.2840 0.1791 0.4290 0.3163
5-shot 0.0821 0.1426 0.0345 0.0502 0.1687 0.2573 0.2303 0.3063 0.0300 0.0492 0.4880 0.4606 0.4450 0.4062 0.4890 0.4509
GPF 1-shot 0.0503 0.0856 0.0027 0.0048 0.1099 0.1702 0.1599 0.2326 0.0072 0.0128 0.3080 0.1952 0.3110 0.1984 0.4220 0.2670
5-shot 0.0839 0.1460 0.0027 0.0047 0.0817 0.1392 0.2014 0.2994 0.0179 0.0310 0.5550 0.5233 0.4690 0.4223 0.5150 0.4934
GPrompt 1-shot 0.0541 0.1102 0.0138 0.0207 0.0797 0.1310 0.1362 0.2073 0.0074 0.0120 0.3540 0.1596 0.2800 0.1519 0.4710 0.3705
5-shot 0.0769 0.1396 0.0157 0.0231 0.1340 0.2166 0.2157 0.3147 0.0287 0.0464 0.5510 0.5098 0.5570 0.5211 0.5130 0.4520
GraphCL 1-shot 0.0603 0.1112 0.0265 0.0398 0.0390 0.0799 0.1655 0.2529 0.0047 0.0077 0.2430 0.1548 0.2980 0.1630 0.4070 0.4130
5-shot 0.0740 0.1368 0.0311 0.0456 0.1416 0.2138 0.2019 0.3075 0.0270 0.0440 0.5610 0.5330 0.4300 0.3683 0.5230 0.5024
OpenGraph 0-shot 0.0921 0.1746 0.0421 0.0639 0.1911 0.2978 0.2370 0.3265 0.0485 0.0748 0.7504 0.7426 0.6097 0.5821 0.6869 0.6537
4.1.4. Baselines
@@
Our empirical evaluation utilizes the following 9 state-of-the-art baseline methods from 4 different research lines. Detailed descriptions for the baselines can be found in Appendix A.2.4.
4.2. Overall Performance Comparison (RQ1)

### v4

@@ 3.1.2 Topology-aware Projection with Arbitary Graphs
To handle the varying dimensionality $|\mathcal{V}|\times|\mathcal{V}|$ of adjacency $\tilde{\textbf{A}}$, OpenGraph applies a projection function $\phi:\mathbb{R}^{|\mathcal{V}|}\rightarrow\mathbb{R}^{d}$ to transform the adjacency into sequence data. A large hidden dimensionality $d$ is used to minimize information loss. Previous research has shown that even random projections with large dimensions can achieve satisfactory performance Zheng et al. (2022). To preserve topology information, we employ fast singular value decomposition (SVD) as the projection $\phi$. SVD is known for its efficiency and effectiveness in adjacency compression Jamali and Ester (2010). Our empirical analysis demonstrates that two iterations of fast SVD effectively preserve topology information with minimal computational overhead. The graph tokenizer performs the following operations to calculate the resulting token sequence:
@@
$\textbf{e}_{v}=\phi(\tilde{\textbf{A}}_{v,:})=\tilde{\textbf{A}}_{v,:}\cdot\text{LN}((\textbf{U}\sqrt{\Lambda}\leavevmode\nobreak\ \|\leavevmode\nobreak\ \textbf{V}\sqrt{\Lambda}))$ (4)
@@
3.2 Scalable Graph Transformer
@@
With the universal topology-aware graph tokens, the subsequent task is to empower our graph model to grasp the complex node-wise dependencies within the global context. Inspired by the success of transformer architectures in modeling complex relationships between instances, OpenGraph utilizes a graph transformer as the backbone. To ensure scalability and effectiveness for large-scale graphs, we introduce the following techniques.
@@
Token Sequence Sampling. For efficiency, we train the graph transformer using sampled token sequences from the current training batch, which contains centric nodes $v_{c_{b}}$, positive nodes $v_{p_{b}}$, and negative nodes $v_{n_{b}}$. The input is as follows:
$(\textbf{e}_{c_{1}}\cdots\textbf{e}_{c_{B}})\leavevmode\nobreak\ \|\leavevmode\nobreak\ (\textbf{e}_{p_{1}}\cdots\textbf{e}_{p_{B}})\leavevmode\nobreak\ \|\leavevmode\nobreak\ (\textbf{e}_{n_{1}}\cdots\textbf{e}_{n_{B}})$ (5)
This approach significantly reduces the sequence length from $|\mathcal{V}|$ to $3\times B$, enabling efficient training for large-scale graphs. Despite using a sub-sequence, the topology-aware embeddings contain local structural information for each node and reflect the overall graph structure. Additionally, this sampling technique emphasizes the current training batch, leading to further training improvements.
Efficient Self-Attention with Anchors. To accelerate the self-attention part of OpenGraph with quadratic complexity, we introduce a step of sampling anchor nodes $v_{a_{s}}$ for $s\in S$, where $S<3B$. This splits the self-attention process into two stages: propagating messages from all nodes to the anchor nodes and then propagating the anchor embeddings to all nodes. This decomposition reduces the complexity from $\mathcal{O}(B^{2}\times d)$ to $\mathcal{O}(B\times S)$, ensuring scalability for large-scale graphs.
3.3 Knowledge Distillation from LLM
@@
Obtaining diverse graph datasets for different domains can be challenging due to factors like privacy issues that restrict access to essential data Zheleva and Getoor (2007). Inspired by the remarkable knowledge and understanding demonstrated by large language models (LLMs), we leverage their power to enhance the generation of diverse graph-structured data. To improve the efficacy of our LLM-augmented graph data for pre-training our model, we have developed an augmentation mechanism. This mechanism enables the LLM-augmented graph data to closely approximate real-world graph characteristics, enhancing the relevance and usefulness of the augmented data.
3.3.1 LLM-based Node Generation
@@
Our first step is to create a node set tailored to the application scenario, characterized by text-based profiles that generate subsequent edges. However, dealing with real-world scenarios poses challenges due to the large scale of the node set. For example, e-commerce platforms may have billions of products, making it challenging for the LLM to efficiently generate a large number of nodes.
@@
To address this challenge, we adopt an iterative strategy of dividing general nodes into sub-categories with finer semantic granularity. For instance, in the case of generating product nodes, we prompt the LLM with a query like "List sub-categories of products on platforms like Amazon." The LLM provides a list of sub-categories such as "clothing" and "electronics." We repeat this iterative division process, refining each sub-category further, until we obtain nodes that resemble real-world instances, such as "women’s clothing," "sweaters," "hooded sweaters," and "white hooded sweaters." Appendix A.2.1 presents details on our prompt template and generation examples.
@@
3.3.2 Edge Sampling using Node Profiles
@@
To generate edges, we use the Gibbs sampling algorithm Gelfand (2000) with the generated node set $\mathcal{V}$. The algorithm starts with a random sample.For instance, in a paper-wise citation network, the initial sample is a node pair $(v_{s_{0}},v_{t_{0}})$. In a person-entity relation scenario like an author-paper network, the initial sample is a binary vector $\textbf{a}^{0}\in\{0,1\}^{|\mathcal{V}|}$. Each element $a_{i}\in\textbf{a}^{0}$ indicates whether there is an interaction between the sampled person and the $i$-th node $v_{i}$. In the case of person-entity relations, the Gibbs algorithm for edge sampling is described in Appendix A.2.3. The key is estimating the probability $p(\mathbf{a}^{t}\oplus v_{t^{\prime}}|\mathbf{a}^{t})$, with $\oplus$ representing setting the $t^{\prime}$-th dimension of $\textbf{a}^{t}$ to 1.
@@
Node-wise Connection Probability Estimation. To estimate the probability $p(\mathbf{a}^{t}\oplus v_{t^{\prime}}|\mathbf{a}^{t})$ of connecting two nodes in our generated graph, we leverage the reasoning capabilities of the LLM. However, directly prompting the LLM for predictions on each edge can be computationally expensive, with $\mathcal{O}(|\mathcal{V}|\times|\mathcal{V}|)$ prompts required. To ensure efficiency, we adopt an alternative approach. We prompt the LLM to generate hidden representations $\mathbf{h}_{i}$ for each node $v_{i}$. Then, we calculate the probability for each edge with dot-product as:
$p(\textbf{a}^{t}\oplus v_{t^{\prime}}|\textbf{a}^{t})=\sum\nolimits_{v_{i}}a^{t}_{i}(\textbf{h}_{i}/\|\textbf{a}^{t}\|_{0})^{\top}\cdot\textbf{h}_{t^{\prime}}$ (6)
@@
By utilizing the text embeddings $\textbf{h}_{i}$ and $\textbf{h}_{t^{\prime}}$ provided by the LLM, we can effectively capture the semantic relations between the respective nodes.
@@
Dynamic Probability Normalization. To ensure that the calculated probability scores fall within a reasonable range like $[0,1]$, our generation algorithm incorporates a dynamic probability normalization approach.It maintains a record of the most recent $T^{\prime}$ estimation values, denoted as $\mathcal{P}=\{p(\textbf{a}^{t}\oplus v_{t^{\prime}}|\textbf{a}^{t})\leavevmode\nobreak\ |\leavevmode\nobreak\ t=-1,\cdots,-T^{\prime}\}$. By calculating the mean ($\mu$) and standard deviation ($\sigma$) of these values, we gain insight into their distribution. New estimations are adjusted using $\mu\pm 2\sigma$ as the bounds, resulting in $\bar{p}=(p-\mu)/(4\sigma)$.
@@
Node Locality Incorporation. To address the limitation of the previous edge sampling algorithm based on semantic similarity, we introduce the concept of locality. Each node is assigned a random locality index, and we consider the difference in locality using an exponential decay function. This results in an adjusted probability calculated as $\hat{p}=\bar{p}\cdot\alpha^{|n_{i}-n_{j}|}$, where $10<\alpha<$ 1. By incorporating locality, we account for the observed patterns in real-world graphs and prevent excessive connections among semantically-related nodes.
3.3.3 Graph Topological Pattern Injection
@@
4 Evaluation
4.1 Experimental Settings
Datasets. We evaluate OpenGraph on two graph learning tasks: link prediction and node classification, using totally 8 real-world datasets. Appendix A.3.1 provides detailed descriptions.
Evaluation Protocol. Following previous works He et al. (2020); Kipf and Welling (2017), we adopt the original train-test data split for the experimental datasets. We pre-train our OpenGraph on generated datasets and conducts zero-shot prediction for the evaluation datasets made of real graph data. As most baselines struggle with cross-dataset transferring, we evaluate them in two few-shot training settings. Please refer to Appendix A.3.2 for more details about our cross-dataset zero-shot setting, few-shot settings, and evaluation metrics.
Implementation Details. We provide detailed information about the implementation of OpenGraph and the baseline methods, as well as the graph generation process, in Appendix A.3.3.
Table 1: Performance of OpenGraph (zero-shot) and baseline methods (one-shot, five-shot) on link prediction (measured by Recall@N for $N=20,40$) and node classification (measured by Accuracy and Macro F1 Score).
@@ 4.1 Experimental Settings
Model shot R-20 R-40 R-20 R-40 R-20 R-40 R-20 R-40 R-20 R-40 Acc F1 Acc F1 Acc F1
MF 1 .0087 .0161 .0261 .0349 .0331 .0604 .1396 .1956 .0034 .0043 .1710 .1563 .1740 .1727 .3470 .3346
5 .0536 .0884 .0412 .0609 .0987 .1584 .2060 .2989 .0196 .0284 .1500 .1422 .1520 .1484 .3540 .3435
MLP 1 .0195 .0336 .0112 .0185 .0548 .1019 .1492 .2048 .0017 .0028 .2300 .1100 .2590 .1993 .4430 .3114
5 .0621 .1038 .0115 .0185 .0851 .1470 .2362 .2563 .0092 .0152 .3930 .3367 .3690 .3032 .5240 .4767
GCN 1 .0279 .0459 .0206 .0321 .0432 .0849 .1760 .2086 .0096 .0160 .3180 .1643 .3200 .2096 .4270 .3296
5 .0705 .1312 .0366 .0513 .1054 .1656 .2127 .2324 .0251 .0408 .5470 .5008 .4910 .4190 .509 .4455
GAT 1 .0580 .1061 .0258 .0372 .0245 .0520 .1615 .2476 .0047 .0079 .2420 .1687 .2810 .2025 .4720 .3657
5 .0711 .1309 .0340 .0505 .1506 .2267 .2002 .2883 .0228 .0392 .585 .5438 .4940 .4441 .5780 .5582
GIN 1 .0530 .1004 .0163 .0247 .0466 .0884 .1541 .2388 .0069 .0114 .3190 .1753 .2820 .1705 .4410 .3064
5 .0735 .1441 .0311 .0458 .1458 .2344 .1926 .2829 .0252 .0418 .5400 .4941 .521 .4696 .5070 .4547
DGI 1 .0315 .0617 .0255 .0385 .0486 .0863 .1868 .2716 .0081 .0142 .3150 .1782 .2840 .1791 .4290 .3163
5 .0821 .1426 .0345 .0502 .1687 .2573 .2303 .3063 .0300 .0492 .4880 .4606 .4450 .4062 .4890 .4509
GPF 1 .0503 .0856 .0027 .0048 .1099 .1702 .1599 .2326 .0072 .0128 .3080 .1952 .3110 .1984 .4220 .2670
5 .0839 .1460 .0027 .0047 .0817 .1392 .2014 .2994 .0179 .0310 .5550 .5233 .4690 .4223 .5150 .4934
GPrompt 1 .0541 .1102 .0138 .0207 .0797 .1310 .1362 .2073 .0074 .0120 .3540 .1596 .2800 .1519 .4710 .3705
5 .0769 .1396 .0157 .0231 .1340 .2166 .2157 .3147 .0287 .0464 .5510 .5098 .5570 .5211 .5130 .4520
GraphCL 1 .0603 .1112 .0265 .0398 .0390 .0799 .1655 .2529 .0047 .0077 .2430 .1548 .2980 .1630 .4070 .4130
5 .0740 .1368 .0311 .0456 .1416 .2138 .2019 .3075 .0270 .0440 .5610 .5330 .4300 .3683 .5230 .5024
Ours 0 .0921 .1746 .0421 .0639 .1911 .2978 .2370 .3265 .0485 .0748 .7504 .7426 .7221 .6801 .6869 .6537
@@
Baselines. Our empirical evaluation utilizes the following 9 state-of-the-art baseline methods from 4 different research lines. Detailed descriptions for the baselines can be found in Appendix A.3.4.
4.2 Overall Performance Comparison (RQ1)